OpenStudy (anonymous):

y+8x= 2x squared +1 solve using complete square

OpenStudy (anonymous):

@freckles

pooja195 (pooja195):

Can you screenshot the question?

OpenStudy (anonymous):

no i can't sorry

pooja195 (pooja195):

I'm not sure what the question is asking.... @zzr0ck3r ??

OpenStudy (anonymous):

y+8x= 2x squared +1

OpenStudy (anonymous):

y+8x= 2x squared +1 solve using complete the square

pooja195 (pooja195):

y+8x= 2x^1?

OpenStudy (anonymous):

y+8x= 2x^2 +1?

pooja195 (pooja195):

is that the question?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

@zzr0ck3r im not sure can u help with this one please? :)

OpenStudy (astrophysics):

\[y=2x^2-8x+1\] now you may go ahead and complete the square

OpenStudy (anonymous):

i got x^2 -8x+32=15 not sure how to factor that

OpenStudy (astrophysics):

\[x^2+ax=0\] we can do the following \[x^2+\frac{ 2ax }{ 2 } + \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2=0\]

OpenStudy (astrophysics):

That's completing the square

OpenStudy (anonymous):

how do we get the x values

OpenStudy (astrophysics):